OpenStudy (rayman5766):

there is a medal for any one who can help me Which shows the expression below in simplified form? (8 × 104) × (7 × 10-5) A. 5.6 × 10-1 B. 5.6 × 100 C. 15 × 10-1 D. 5.6 × 101

OpenStudy (rayman5766):

@logan13 hey bud can you help me please

OpenStudy (rayman5766):

@leonardo0430 hey can you help me just for a sec

OpenStudy (anonymous):

(8)(104)((7)(10)−5) =((8)(104))((7)(10)+−5) =((8)(104))((7)(10))+((8)(104))(−5) =58240−4160 =54080

OpenStudy (rayman5766):

say what!